# An Item is Worth a Prompt: Versatile Image Editing with Disentangled Control

### v2

@@ 3.4 Editing with Item-Prompt Freestyle
Image-based Item Editing. Given two images, to replace an item in the source image with another one in the reference image, all item prompts in both images should be injected into the same model. Then the item replacement can be realized by the corresponding prompt replacement Note that this setting is more general compared to classical image-based editing scenarios where the reference is an entire image instead of an item. This is particularly useful when it is hard to find a solo focus reference.
@@
4 Experiments
Figure 5: D-Edit enables text-guided editing in complex images. D-Edit enables the user to select any of the items from segmentation and replace it with the target text prompt.
4.1 Text-Guided Editing
In Fig. 6(b), we show the results of item personalization where we generate the image using certain item prompt(s), without using the item-prompt association. This differs from Dreambooth-style personalization in that Dreambooth lacks the capability for item-level customization unless the item is cropped and focused upon, which is usually hard in an image including multiple images. Besides, it requires more images (3-5) with captions for personalization, while our method takes one image without captions. Qualitative results show that the learned item(s) can be combined with text to generate personalized images, and the text prompt can be used to personalize the background, number, and position of the item.
Figure 6: Besides text replacement, the learned prompt (denoted as [v]) can be combined with text words to achieve refinement/editing of the target item. (a) Augment an item prompt with text words while keeping other prompts unchanged for editing. (b) Generating the entire image with certain item prompt(s) augmented with text words for personalization.
4.2 Image-Guided Editing
For image-guided editing, the user can select one item from the reference image and use it to replace one item in the target image. We compare the editing results with baselines including Anydoors [3], Paint-by-Example [26] and TF-ICON [8] when the reference image mainly consists of a single item. As shown in Fig. 7, Paint-by-example can naturally inpaint reference items into the target scene but it falls short in keeping the identity of the reference item (the face and bird example). Anydoors can retain more relevant details from the reference image, yet it may also incorporate undesirable elements in reference, resulting in a less harmonious blend with the target image. For example, the car’s original orientation is preserved, causing it to appear out of the parking spot in the target image. Besides, it cannot preserve the face details as in the example. Compared to these methods, D-Edit is capable of seamlessly composing objects into the target while maintaining their identities.
Figure 8: Image-guided editing using D-Edit. Users can select any target item in the input image and replace it with another item from the same or different reference images.
Figure 9: Different types of mask-based editing: (a) moving/swapping items; (b) reshaping a (part of) an item; (c) resizing an item.
4.3 Mask-Based Editing and Item Removal
D-Edit also enables removing items by deleting the mask-item-prompt pairs. In Fig. 11, by deleting items from the scene image one by one, the resulting blank region will be re-partitioned to nearby masks and join the corresponding item-prompt pair. D-Edit will then use such new associations to imagine the blank regions and therefore lead to reasonable filling results.
Figure 11: Results of removing items one by one from the original image.
5 Conclusion
In this work, we propose D-Edit as a versatile image editing framework based on diffusion models. D-Edit segments the given image into multiple items, each of which is assigned a prompt to control its representation in the prompt space. The image-prompt cross-attention is disentangled to a group of item-prompt interactions. Each prompt is constrained to exclusively interact with the item it governs through isolated cross-attention, therefore disentangling the cross-attention control pipelines. Item-prompt associations are built up by fintuning the diffusion model which learns to reconstruct the original image using the given set of item prompts. D-Edit accomplishes image editing by allowing users to freely change or establish associations between prompts and items, and we highlight four types of image editing operations including text-based, image-based, mask-based editing, and item removal. We showcase the quality and versatility of the editing results across a diverse range of collected images through both qualitative and quantitative evaluations. Compared to baselines, D-Edit can better preserve the information of the original image while following versatile guidance.

### v3

@@ Appendix A Different types of editing
Image-based Item Editing. Given two images, to replace an item in the source image with another one in the reference image, all item prompts in both images should be injected into the same model. Then the item replacement can be realized by the corresponding prompt replacement. Note that this setting is more general compared to classical image-based editing scenarios where the reference is an entire image instead of an item. This is particularly useful when it is hard to find a solo focus reference.
@@
Appendix B Prompt Interpolation
Instead of replacing the original item prompt with the target one (image/text guided), we can also interpolate between the two and get the transition results. As shown in Fig. 13, we interpolate the embeddings after the text encoder and derive the mixed embedding using $c=\text{alpha}\cdot c_{\text{guide}}+(1-\text{alpha})\cdot c_{\text{orig}}$.
Figure 13: Mixing of the original and guidance images by linearly interpolating the embeddings of the two.
Appendix C More Text-Based Editing
Figure 14: More text-editing results of replacing one item with target prompts.
Appendix D More Image-Based Editing
As marked in Fig. 15, several failure examples occur: (1) Color is harder to learn than shape: in the Lamborghini example $a$, the shape is preserved while the yellow color is lost. (2) Item should have clear meanings: in the man-in-car example $b$, because the man picture through the window is blurred and DPM cannot understand, the inpainting results will be unnatural. (3) Out-of-domain image is hard to learn: In the celebrity face example $c$, if the celebrity face is not natural or highly Photoshopped, it will be hard for DPMs to reconstruct it and therefore lead to weird face results.
Figure 15: More image-guided editing results of replacing one item with the reference item. Failure examples are marked by red letters.